supdasup:

An ice cream cone has a diameter of 3 in. and a height of 5 in. Find the volume to see how much ice cream will fit inside the cone. Use 3.14 for 𝜋. Round your answer to the nearest whole number.

supdasup:

@563blackghost

supdasup:

@Hero

supdasup:

@Ultrilliam

Ultrilliam:

@Shadow When your on @563blackghost

Hero:

@supdasup, do you remember the formula for volume of a cone?

supdasup:

No i dont

supdasup:

Mainly why i need help

Hero:

The volume of a cone \(V\) with radius \(r\) and height \(h\) is \(V = \dfrac{1}{3}\pi r^2h\)

Hero:

There's your formula. And you are given the values for radius and height. Remember to use 3.14 for \(\pi\). Let me know what you get for \(V\) when you're finished calculating.

supdasup:

so 8 ?

Hero:

May I ask how you got 8? Can you show your work here please?

supdasup:

by adding the inches lol

Hero:

No, you don't add. You multiply.

supdasup:

wait, is it 15 then?

Hero:

You're definitely guessing.

supdasup:

yeah

supdasup:

just give me the answer and im good

supdasup:

I am guessing becuz i have no idea

Hero:

Do you not know how to multiply? That's all you have to do here.

supdasup:

multiply what? the inches?

Hero:

\(\dfrac{1}{3} \times 3.14 \times (2.5)^2 \times (5)\)

supdasup:

sorry i had to help my dad

supdasup:

so, if just multiplying that, then 32?

Hero:

Close enough although there should be some decimals as well. You should not round off the decimals.